# Policy Search, Retrieval, and Composition via Task Similarity in Collaborative Agentic Systems

### v2

@@ 3.4 Knowledge Composition and Fine Tuning
Masks received from other agents meeting the two selection criteria above have been pretrained on tasks that are likely related to, or share similarities with, the agent’s current task. As previously shown in Ben-Iwhiwhu et al. (2023), a linear combination of masks, weighted with trainable parameters $\beta$, can lead to beneficial policy search in RL. MOSAIC exploits this idea by combining policies that have been trained on the local agent, plus those received from other agents. In this study, the $\beta$ values are computed via softmax over a set of beta parameters optimized in log-space, $\tilde{\beta}\in\mathbb{R}^{1+|P_{c}|}$, where $P_{c}$ is the number of masks received. During the forward pass, the agent constructs the linearly combined, binarized mask,
@@
where $\phi_{\tau}$ is the agent’s own task mask, and $\phi_{k}$ is the $k$-th peer mask. The parameters $\beta_{\tau}$ and $\beta_{k}$ are the softmax-normalized values derived from $\tilde{\beta}$, for $\phi_{\tau}$ and $\phi_{k}$, respectively. The resulting binary mask, $\phi_{\tau}^{\text{lc}}$, is then used to modulate the backbone parameters, yielding the final policy $\pi_{\tau}$. As training continues on the local task, the agent updates the $\beta$ parameters and $\phi_{\tau}$ via backpropagation (keeping all the $\phi_{k}$ masks fixed) thus fine tuning the overall resulting policy determined by $\phi_{\tau}^{\text{lc}}$.
@@
The trainable $\beta$ parameters enable gradient descent to determine which policies are most useful for the current task. The initial weighting as the masks are received is given by
$\beta_{\tau}=0.5+0.5r,\quad\beta_{k}=\frac{0.5(1-r)}{|P_{c+1}|}\quad,$ (8)
where $r\in[0,1]$ is the agent’s normalized return from the last iteration and $|P_{c+1}|$ is the number of received peer masks. This scheme biases low-performing agents toward external knowledge and high-performing agents toward their own policies. Ablation studies show that such an initialization provides a strong advantage, even though $\beta$ is later tuned by gradient descent.
@@
The consolidation collapses multiple masks into one without affecting the policy, maintaining memory scalability.
@@ 4 Experiments
MOSAIC is evaluated on three sparse-reward reinforcement learning benchmarks. The CT-graph (Soltoggio et al. 2019) and MiniHack MultiRoom (Samvelyan et al. 2021) are used to assess the advantage of communicating agents versus isolated agents. The MiniGrid Crossing (Chevalier-Boisvert, Willems, and Pal 2018) is used to assess the performance of MOSAIC against the chosen baselines. Each benchmark includes tasks with similar, dissimilar and interfering tasks to assess whether MOSAIC can identify and leverage the available task similarities, while avoiding interfering policies. Each agent is assigned a unique task and trained in parallel with other agents. PPO (Schulman et al. 2017) was used across all experiments, with an FCN for MiniGrid and the CT-graph, and a CNN for MiniHack. Individual results are shown in Appendix D, Figures 14, 15, and 16. Details on computing infrastructure, hyperparameters, architectures, and libraries are reported in the Appendix E, Tables 3, 4, 5, and 6.
@@ 4.1 Image Sequence Learning
The image sequence learning (ISL) benchmark, implemented with the Configurable Tree Graph (CT-Graph) environment (Soltoggio et al. 2019), consists of procedurally generated tree navigation problems where each node is an RL state encoded as an image. We define a curriculum of 28 tasks organized into four independent image sets, each containing seven related tasks of increasing tree depth (2–8). This configuration produces four distinct and unrelated task groups with internal hierarchies of difficulty, from easy (depth 2) to very hard (depth 8). Sparse rewards and exponential branching make the benchmark challenging, with reward probabilities as low as $\approx 7.74\times 10^{-9}$ for depth-8 tasks (Soltoggio et al. 2019).
@@ 4.3 MiniGrid Crossing
The MiniGrid benchmark is a sparse-reward grid-world navigation task with symbolic observations. MOSAIC is evaluated on 14 tasks spanning seven SimpleCrossing and seven LavaCrossing variants, which differ in layout, object placement, and room structure. Summed return curves for all methods are shown in Figure 3, with final and intermediate metrics in Table 1. MOSAIC is compared against the following baselines. Multi-Task PPO (MTPPO) uses a shared backbone without modularity or interference mitigation. Multi-DQN (MDQN) (D’Eramo et al. 2020) attaches task-specific Q-heads to a shared encoder but has no mechanism to avoid interference. PCGrad+MoE (Yu et al. 2020) combines a shared encoder, expert subnetworks, and gradient projection to resolve conflicts. Mixture of Orthogonal Experts (MOORE) (Hendawy, Peters, and D’Eramo 2024) promotes expert diversity via orthogonalization with learned gating to reduce overlap.
@@ 4.4 How Similarity Helps Targeted Policy Transfers
Figure 4 illustrates the average cosine similarity in the CT-graph benchmark between task embeddings and the converged consolidation parameters $\hat{\beta}^{i}_{j}$ (see Equation (9)), which provides an indication of the influence of policies derived from task $j$ when learning task $i$.
@@
Figure 5 presents the performance of agents in the CT-graph grouped by difficulty levels. Easier levels are learned first and harder levels later. Interestingly, the comparison with the isolated agents (bottom graph) suggests that the policies of easier tasks are progressively shared with agent solving harder tasks, enabling those agents to find successful policies that the same RL algorithm fails to discover when learning in isolation. This analysis helps explain the 170.8% performance gain of communicating agents over isolated agents (Figure 2(A)): agents tackling harder tasks retrieve, combine, and refine policies from simpler tasks to solve complex problems.
Figure 4: Pairwise cosine similarity and $\hat{\beta}$ statistics in the image sequence learning benchmark. (A) Cosine similarity matrix. (B) $\hat{\beta}$ values indicating policy use per task. (C) Cosine similarities clustered using WPGMA, shown as a dendrogram. (D) Cosine similarity matrix reordered by clustering. (E) Clustered $\hat{\beta}$ values show that similar tasks exhibit the highest policy reuse. Annotated heatmaps of individual clusters are in Appendix D.1, Figure 18.
Figure 5: Performance grouped by task complexity, i.e., seven different levels, corresponding to different graph’s depths, in the image sequence learning problem. Average performance for 20 runs (four tasks and five seeds/task) with 95% confidence intervals. Communicating MOSAIC agents (top graph) are compared with isolated MOSAIC agents (bottom graph). Communicating agents solve tasks progressively, from the simplest to the hardest tasks. Isolated agents only manage to solve the two simplest tasks, and partially the third, but fail on the four most complex tasks.
@@ 4.5 Ablation Studies
Ablation studies were conducted to assess the effect on performance of MOSAIC’s knowledge selection criteria and reward-guided initialization (RGI) on performance. Figure 6 compares three ablated variants of MOSAIC: without cosine similarity-based selection ($\neg$ crit. 1), without performance-based selection ($\neg$ crit. 2), and without reward-guided weight initialization ($\neg$ RGI). MOSAIC-NoComm is included as a reference baseline.
Ablation of $\neg$ criterion 1 reaches a total return of 20.1, whereas MOSAIC reaches 26.0, outperforming it by a factor of 1.29. We speculate that this gap could widen further in settings in which the number of unrelated tasks grows, i.e., where selecting the most relevant knowledge becomes critical.
@@
The third ablation, $\neg$ RGI removes the reward-guided initialization and instead uses a fixed weighting of 0.5 for the local policy and 0.5 for the external policies. This experiment performs comparably with MOSAIC; however, the average return shows periodic instability, characterized by sharp performance drops coinciding with communication events, resembling a sawtooth pattern. This behavior suggests that reward-guided initialization is a useful approach to maintaining performance stability when integrating external knowledge.
Further ablation studies are shown in Appendix C to test the impact of query frequency on the performance of MOSAIC, and the impacts of the number of samples and size of the reference distribution on embedding accuracy and performance.
Figure 6: Figure 6: Ablation studies showing average performance over 140 runs (28 tasks, five seeds per task), with 95% confidence intervals. Variants tested include MOSAIC with $\neg$ (Criterion 1) (no similarity criterion), $\neg$ (Criterion 2) (no reward criterion), or $\neg$ RGI (no reward-guided initialization). MOSAIC-NoComm is included as a baseline. Removing any of these components significantly degrades performance, highlighting their combined importance.
@@ 5 Discussion
The results show that MOSAIC’s principles are effective across diverse settings. Agents achieved faster and more successful learning dynamics than isolated agents. In particular, transferring masks from some tasks accelerated learning in other tasks, consistent with findings in centralized lifelong learning (Ben-Iwhiwhu et al. 2023). These gains were most evident in reward-sparse and hierarchical environments, where agents learning simpler tasks supported those tackling harder ones (Florensa et al. 2018; Nachum et al. 2018; Uchendu et al. 2023).
The ablation studies highlight the importance of the selection criteria (cosine similarity for relevance and reward-based selection for utility) as well as reward-based weighting when integrating new policies. Without these criteria, knowledge sharing could degrade the performance (Isele and Cosgun 2018) and thus emphasize the importance for scalability when deploying large numbers of agents on diverse tasks (Foerster et al. 2016). The effect of reward-guided initialization suggests that careful policing weighting is important to avoid performance disruption before fine tuning (Amodei et al. 2016).
It is worth noting that MOSAIC’s design takes inspiration from lifelong learning algorithms. Thus, policy composition can be easily extended to previously learned policies of past tasks. Such an extension could deal with continuous streams of evolving tasks (Serra et al. 2018), supporting scalable and open-ended lifelong learning (Leibo et al. 2019; Stanley 2019).
@@
MOSAIC relies on raw episodic reward as a selection and weighting signal, which may limit generalization of the method to environments with differing reward functions. Reward-normalized scores or task-progress metrics could improve environment-agnostic reuse (Jaques et al. 2019). Scalability is also affected by the mesh topology; adaptive or sparse topologies would better mitigate bandwidth constraints (Kaimakamidis, Mademlis, and Pitas 2024; Marino et al. 2025; Tang et al. 2024). The policy composition used in MOSAIC is limited to only positive combinations. Other composition or mixture approaches could be tested (He et al. 2022; Wu, Huang, and Wei 2024). Alternative modular skill representations could be used for applications to larger neural networks, e.g., LoRA-like modules and adapters for transformer architectures (He et al. 2022; Wu, Huang, and Wei 2024; Han et al. 2021; Qiu et al. 2020). In all such cases, the requirement for a shared backbone or foundation model may reduce the pool of agents able to share knowledge and limit long-term development of progressively more complex skills (Houlsby et al. 2019).
Communication and selection mechanisms in MOSAIC were introduced as a proof-of-concept to test whether selective transfer is essential for agentic knowledge reuse (DARPA/MTO 2021; Soltoggio et al. 2024). Allowing policies to learn when and how to communicate could further increase autonomy (Foerster et al. 2016), but dynamic communication policies raise safety concerns. Agents could learn uncooperative strategies, exploit collective knowledge without contributing, or share harmful policies. Such risks mirror challenges in Game Theory (Myerson 2013) and require safegaurds (Hammond et al. 2025; Bengio et al. 2025). MOSAIC-like systems may also face adversarial attacks exploiting the agent-to-agent nature through model poisoning or policy extraction (Vassilev et al. 2024).
@@ 6 Conclusion
This work introduced Modular Sharing and Composition in Collective Learning (MOSAIC), where agents search, retrieve, and compose policies obtained from peers learning other tasks. The collective was shown to significantly outperform isolated agents by selecting and fine-tuning task-aligned knowledge. Results show that policy sharing is most effective when implemented selectively with policy weighting and fine-tuning. Collective learning also results in implicit curricula in which agents solving simpler tasks accelerate progress on harder ones, enabling faster learning of complex tasks.
MOSAIC also improves interpretability, as composed policies are traceable to their sources, and the approach could extend to domains such as language, perception, and control. However, autonomous knowledge sharing carries risks: flawed or misaligned policies may spread rapidly. Future work must ensure that transfers remain safe and aligned with agent objectives. By leveraging task similarities through selective knowledge reuse, MOSAIC could supports scalable, adaptive, and collaborative agentic AI in real-world applications.
Appendix A Extended Related Work

### v3

@@ 3.4 Knowledge Composition and Fine Tuning
Masks received from other agents meeting the two selection criteria above have been pretrained on tasks that are likely related to or share similarities with the agent’s current task. As previously shown in Ben-Iwhiwhu et al. (2023), a linear combination of masks, weighted with trainable parameters $\beta$, can lead to beneficial policy search in RL. MOSAIC exploits this idea by combining policies that have been trained on the local agent, plus those acquired from other agents. In this study, the $\beta$ values are computed using softmax on a set of beta parameters, optimized in log-space, $\tilde{\beta}\in\mathbb{R}^{1+|P_{c}|}$, where $P_{c}$ is the number of masks acquired during a communication event $c$. During the forward pass, the agent constructs the linearly combined, binarized mask,
@@
where $\phi_{\tau}$ is the agent’s own task mask, and $\phi_{k}$ is the $k$-th peer mask. The parameters $\beta_{\tau}$ and $\beta_{k}$ are the softmax-normalized values derived from $\tilde{\beta}$, for $\phi_{\tau}$ and $\phi_{k}$, respectively. $|P_{c+1}|$ is the number of masks acquired in the next communication event. The resulting binary mask, $\phi_{\tau}^{\text{lc}}$, is then used to modulate the backbone parameters, which gives the final policy $\pi_{\tau}$. As training continues on the local task, the agent updates the $\beta$ parameters and the real value $\phi_{\tau}$ via backpropagation (keeping all $\phi_{k}$ masks fixed) thus fine tuning the overall resulting policy determined by $\phi_{\tau}^{\text{lc}}$.
@@
The trainable parameters, $\beta$, enable gradient descent to determine which policies are most useful for the current task. The initial weighting as the masks are received is given by,
$\beta_{\tau}=0.5+0.5\overline{r},\quad\beta_{k}=\frac{0.5(1-\overline{r})}{|P_{c+1}|}\quad,$ (8)
where $\overline{r}\in[0,1]$ is the agent’s normalized return from the last iteration. This scheme biases low-performing agents toward external knowledge and high-performing agents toward their own policies. Ablation studies show that such an initialization provides a strong advantage, even though $\beta$ is later tuned by gradient descent.
@@
The consolidation collapses multiple masks into one without affecting the policy, managing memory scalability.
@@ 4 Experiments
MOSAIC is evaluated on three sparse-reward reinforcement learning benchmarks. The CT-graph (Soltoggio et al. 2023) and MiniHack MultiRoom (Samvelyan et al. 2021) are used to assess the advantage of communicating agents versus isolated agents. The MiniGrid Crossing (Chevalier-Boisvert, Willems, and Pal 2018) is used to assess the performance of MOSAIC against the chosen baselines. Each benchmark includes tasks with similar, dissimilar, and interfering tasks to assess whether MOSAIC can identify and leverage the available task similarities, while avoiding interfering policies. Each agent is assigned a unique task and trained in parallel with other agents. PPO (Schulman et al. 2017) was used in all experiments, with an FCN for MiniGrid and CT-graph, and a CNN for MiniHack. Individual results are shown in Appendix E, Figures 14, 15, and 16. Details on computing infrastructure, hyperparameters, architectures, and libraries are reported in the Appendix F, Tables 5, 6, 8, and 7. Appendix A, Table 3 reports significance testing.
@@ 4.1 Image Sequence Learning
The image sequence learning (ISL) benchmark, implemented with the Configurable Tree Graph (CT-Graph) environment (Soltoggio et al. 2023), consists of procedurally generated tree navigation problems where each node is an RL state encoded as an image. We define a curriculum of 28 tasks organized into four independent image sets, each containing seven related tasks of increasing tree depth (2–8). This configuration produces four distinct and unrelated task groups with internal hierarchies of difficulty, from easy (depth 2) to very hard (depth 8). Sparse rewards and exponential branching make the benchmark challenging, with reward probabilities as low as $\approx 7.74\times 10^{-9}$ for depth-8 tasks (Soltoggio et al. 2023).
@@ 4.3 MiniGrid Crossing
The MiniGrid benchmark is a sparse-reward grid-world navigation task with symbolic observations. MOSAIC is evaluated on 14 tasks spanning seven SimpleCrossing and seven LavaCrossing variants, which differ in layout, object placement, and room structure. Total return curves for all methods are shown in Figure 3, with final and intermediate metrics in Table 1. MOSAIC is compared against the following baselines: Multi-Task PPO (MTPPO) uses a shared backbone without modularity or interference mitigation. Multi-DQN (MDQN) (D’Eramo et al. 2020) attaches task-specific Q-heads to a shared encoder but has no mechanism to avoid interference. PCGrad+MoE (Yu et al. 2020) combines a shared encoder, expert subnetworks, and gradient projection to resolve conflicts. Mixture of Orthogonal Experts (MOORE) (Hendawy, Peters, and D’Eramo 2024) promotes expert diversity via orthogonalization with learned gating to reduce overlap.
@@ 4.4 How Similarity Helps Targeted Policy Transfers
Figure 4 illustrates the average cosine similarity in the CT-graph benchmark between task embeddings and the converged consolidation parameters $\hat{\beta}^{i}_{j}$ (see Eq. (9)), which provides an indication of the influence of policies derived from task $j$ when learning task $i$.
@@
Figure 5 presents the performance of agents in the CT-graph grouped by difficulty levels. Interestingly, comparing with isolated learning (bottom graph) suggests that the policies of easier tasks are progressively shared with agents solving harder tasks. This coordination enabling those agents to find successful policies that the same RL algorithm fails to discover when learning in isolation. This analysis helps explain the 170.8% performance gain of communicating agents over isolated agents (Figure 2(A)): agents tackling harder tasks retrieve, combine, and refine policies from simpler tasks to solve complex problems.
Figure 4: Pairwise cosine similarity and $\hat{\beta}$ statistics in the image sequence learning benchmark. (A) Cosine similarity matrix. (B) $\hat{\beta}$ values indicating policy use per task. (C) Cosine similarities clustered using WPGMA, shown as a dendrogram. (D) Cosine similarity matrix reordered by clustering. (E) Clustered $\hat{\beta}$ values show that similar tasks exhibit the highest policy reuse. Annotated heatmaps of individual clusters are in Appendix E.1, Figure 17.
Figure 5: Performance grouped by task complexity, i.e., seven different levels, corresponding to different graph depths, in the image sequence learning problem. Average performance for 20 runs (four tasks and five seeds/task) with 95% confidence intervals. Communicating MOSAIC agents (top graph) are compared with isolated MOSAIC agents (bottom graph). Communicating agents solve tasks progressively, from the simplest to the hardest tasks. Isolated agents only manage to solve the two simplest tasks, and partially the third, but fail on the four most complex tasks.
@@ 4.5 Ablation Studies
Ablation studies were conducted to assess the effect on performance of MOSAIC’s knowledge selection criteria and reward-guided initialization (RGI) on performance. Figure 6 compares three ablated variants of MOSAIC: without cosine similarity-based selection ($\neg$ Criterion 1), without performance-based selection ($\neg$ Criterion 2), and without reward-guided weight initialization ($\neg$ RGI). MOSAIC-NoComm is included as a reference baseline.
$\neg$ Criterion 1 reaches a total return of 20.1, whereas MOSAIC reaches 26.0, outperforming it by a factor of 1.29. We speculate that this gap could widen further in settings in which the number of unrelated tasks grows, i.e., where selecting the most relevant knowledge becomes critical.
@@
$\neg$ RGI removes reward-guided initialization and instead uses a fixed weighting of 0.5 for the local policy and 0.5 for external policies. This experiment performs comparably with MOSAIC; however, the average return shows periodic instability, characterized by sharp performance drops coinciding with communication events, resembling a sawtooth pattern. This behavior suggests that RGI is crucial to learning stability when integrating external masks.
Further ablation studies are shown in Appendix D to test the impact of query frequency on the performance of MOSAIC, and the impacts of the number of samples and size of the reference distribution on embedding accuracy and performance.
Figure 6: Ablation studies showing average performance over 140 runs (28 tasks, five seeds per task), with 95% confidence intervals. Variants tested include MOSAIC with no similarity criterion ($\neg$ Criterion 1), no reward criterion ($\neg$ Criterion 2), or no reward-guided initialization ($\neg$ RGI). MOSAIC-NoComm is included as a baseline. Removing any of these components significantly degrades performance, highlighting their combined importance.
@@ 5 Discussion
The results show that MOSAIC’s principles are effective across diverse settings. Agents achieved faster and more successful learning dynamics than isolated agents. In particular, transferring masks from some tasks accelerated learning in other tasks, consistent with findings in centralized lifelong learning (Ben-Iwhiwhu et al. 2023). These gains were most evident in reward-sparse and hierarchical environments, where agents that learned simpler tasks supported those that learned harder ones (Florensa et al. 2018; Nachum et al. 2018).
The ablation study highlights the importance of selection based on relevance and utility and of reward-adjusted weighting when integrating new policies. Naïve sharing overwrites stable learning (Isele and Cosgun 2018), deteriorating performance. These results emphasize the importance of selection when the available tasks and curriculum are unknown (Foerster et al. 2016). The effect of reward-guided initialization suggests that careful policy weighting is important to avoid performance disruption before fine tuning.
It is worth noting that MOSAIC’s design takes inspiration from lifelong learning algorithms. Thus, policy composition can easily be extended to previously learned policies of past tasks. Such an extension could deal with continuous streams of evolving tasks (Serra et al. 2018), supporting scalable and open-ended lifelong learning.
@@
Although the experiments were conducted in simulation, MOSAIC is designed for decentralized, bandwidth-limited deployments (e.g., robot fleets, on-device perception). Agents exchange only compact embeddings and mask scores while inference stays on-device. Key deployment considerations are cross-site normalization, private/authenticated exchange, and tuning query frequency and size to fit bandwidth.
MOSAIC relies on raw per-iteration reward as a selection and weighting signal, which limits generalization to environments with differing reward functions. Reward-normalized scores or task-progress metrics could improve environment-agnostic reuse. Adaptive or sparse topologies would better mitigate bandwidth scalability constraints (Tang et al. 2024). The policy composition used in MOSAIC is limited to positive combinations only. Other composition or mixture approaches could be tested (He et al. 2022). Alternative modular skill representations could be used for applications to larger neural networks, e.g., LoRA-like modules and adapters for transformer architectures (Hu et al. 2022; He et al. 2022). In all such cases, the requirement for a shared backbone or foundation model can reduce the pool of agents able to share knowledge and limit long-term development of progressively more complex skills (Houlsby et al. 2019).
Communication and selection mechanisms in MOSAIC were introduced as a proof-of-concept to test whether selective transfer is essential for agentic knowledge reuse (DARPA/MTO 2021; Soltoggio et al. 2024). Allowing policies to learn when and how to communicate could further increase autonomy (Foerster et al. 2016), but dynamic communication policies raise safety concerns. Agents could learn uncooperative strategies, exploit collective knowledge without contributing, or share harmful policies. Such risks mirror classic game-theoretic challenges and require safeguards. MOSAIC-like systems may also face adversarial attacks exploiting the agent-to-agent nature through model poisoning or policy extraction.
@@ 6 Conclusion
This work introduced Modular Sharing and Composition in Collective Learning (MOSAIC), where agents search, retrieve, and compose policies obtained from peers learning other tasks. The collective was shown to significantly outperform isolated agents by selecting and fine-tuning task-aligned knowledge. Results show that policy sharing is most effective when implemented selectively, with policy weighting and fine-tuning. Selection also results in the discovery of implicit curricula where simpler tasks help agents learn complex ones faster.
MOSAIC improves interpretability, as composed policies are traceable to their sources and the approach could extend to domains such as language, perception, and control. However, autonomous knowledge sharing carries risks: flawed or misaligned policies may spread rapidly. Future work must ensure that transfers remain safe and aligned with agent objectives. Using selective and modular reuse, MOSAIC can support scalable, adaptive, and collaborative agentic AI in real-world applications.
Acknowledgements
This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under contract No. HR001121901 (Shared Experience Lifelong Learning) and the Industrial Robots-as-a-Service (IRaaS) project funded by the EPSRC (EP/V050966/1).
Appendix A Significance Testing
Deep reinforcement learning experiments exhibit substantial variability across random seeds and environment stochasticity (Henderson et al. 2018). To improve the robustness and interpretability of the results, we follow the recommendations of (Colas, Sigaud, and Oudeyer 2018) and report both confidence intervals and formal significance tests on the results.
For each benchmark, each algorithm is evaluated on multiple agents. For every agent, 5 independent algorithm seeds and 5 environment seeds. A difference test is computed using the Welch t-test and bootstrap confidence interval (BCI). Table 3 reports the results of this test for the results of each benchmark.
CT-graph ISL
Method Final Perf (mean $\pm$ 95% CI) $\Delta$ (MOSAIC - Method) Welch $p$ (vs MOSAIC) 95% BCI of $\Delta$
MOSAIC 0.91 [0.87, 0.95] - - -
[…]
MiniGrid Crossing
Method Final Perf (mean $\pm$ 95% CI) $\Delta$ (MOSAIC - Method) Welch $p$ (vs MOSAIC) 95% BCI of $\Delta$
MOSAIC 0.83 [0.75, 0.90] - - -
[…]
MiniHack MultiRoom
Method Final Perf (mean $\pm$ 95% CI) $\Delta$ (MOSAIC - Method) Welch $p$ (vs MOSAIC) 95% BCI of $\Delta$
MOSAIC 0.64 [0.54, 0.74] - - -
[…]
Table 3: Total evaluation performance and significance testing CT-graph, MiniGrid and MiniHack benchmarks. $\Delta$ denotes MOSAIC - Method. Rows highlighted in green indicate methods for which MOSAIC is significantly better (Welch’s $t$-test $p<0.05$ and 95% bootstrap CI of $\Delta$ excludes 0).
Appendix B Extended Related Work